# First Author Affiliation / Address line 1 Affiliation / Address line 2

### v1

@@ Abstract
In this paper, we introduce EasyEdit2, a framework designed to enable plug-and-play adjustability for controlling Large Language Model (LLM) behaviors. EasyEdit2 supports a wide range of test-time interventions, including safety, sentiment, personality, reasoning patterns, factuality, and language features. Unlike its predecessor, EasyEdit2 features a new architecture specifically designed for seamless model steering. It comprises key modules such as the steering vector generator and the steering vector applier, which enable automatic generation and application of steering vectors to influence the model’s behavior without modifying its parameters. One of the main advantages of EasyEdit2 is its ease of use—users do not need extensive technical knowledge. With just a single example, they can effectively guide and adjust the model’s responses, making precise control both accessible and efficient. Empirically, we report model steering performance across different LLMs, demonstrating the effectiveness of these techniques. We have released the source code on GitHub [^1] https://github.com/zjunlp/EasyEdit along with a demonstration notebook. In addition, we provide an online system for real-time model steering, and a demo video [^2] https://zjunlp.github.io/project/EasyEdit2/video for a quick introduction.
@@ 1 Introduction
To this end, we introduce EasyEdit2—a new, easy-to-use steering framework for editing LLMs. Building on the foundation of the legacy EasyEdit Yao et al. (2023); Wang et al. (2024b); Zhang et al. (2024), EasyEdit2 features an entirely new architecture designed to enhance plug-and-play capabilities and improve adjustability when steering LLMs. Currently, a variety of steering methods—including prompt-based steering, activation-based interventions  (Turner et al., 2023; Rimsky et al., 2024; Wang et al., 2024c; Hartvigsen et al., 2023; Scialanga et al., 2025), decoding-based control—exist, yet they remain fragmented and require custom implementations and significant expertise. Thus, we develop the steering vector generator module and the steering vector applier module, to automatically generate steering vectors and use these vectors for intervention (if employing prompt-based steering, generating a steering vector is unnecessary). We also develop a steering vector library to facilitate users in reusing existing steering vectors. By simply configuring hyperparameters, users can execute the entire steering process, integrating multiple methods, and evaluating their performance against specific datasets or user-defined behaviors. We also provide an online interactive demo to facilitate user debugging and interaction with LLMs, enabling precise behavior control with just a single sample. To further assist users, our framework is released under the MIT License, ensuring open access and flexibility for use, modification, and distribution.
We hope that the open-sourcing of the EasyEdit2 can contribute to advancements in natural language processing, machine learning, and research on safety, personalization, and reasoning.
@@
To help users better understand and apply EasyEdit2, we propose the three lines grounded in empirical observations and ethical considerations:
(1) Understanding Strengths and Limitations: Prompt-based approaches often achieve strong performance with minimal setup (Wu et al., 2025), but offer limited control over intervention strength. Activation-based methods provide finer control via a scaling coefficient, though this does not consistently improve performance. Please objectively consider the maturity and limitations of the technology Da Silva et al. (2025); Im and Li (2025).
(2) Empirical Guidelines for Hyperparameter Selection: Middle-to-late layers generally perform better for layer selection. Activation-based approaches rely on a scaling coefficient to control strength, but increasing it does not consistently improve outcomes and may lead to multi-peak or unstable behaviors. This could stem from competing objectives within a single steering vector or deeper nonlinearity in activation space. We encourage further study on the effects of scaling and the mechanism of activation-based steering.
(3) Ensuring Responsible Use: Due to its potential for misuse, any steering conducted with EasyEdit2 should be guided by AI safety measures Chen et al. (2024); Youssef et al. (2025).
@@ 2 Background
Inference-time steering modifies model behavior during inference through prompt-based (Wu et al., 2025), activation-based (Zou et al., 2023; Stolfo et al., 2024; Bartoszcze et al., 2025; Wehner et al., 2025), and decoding-based methods (Liang et al., 2024). Compared to parameter fine-tuning methods Han et al. (2024b), inference-time intervention offers several key advantages: (1) Pluggability—steering methods can be seamlessly applied or removed without changing model weights, whether through activation modification, prompt-based guidance, or decoding adjustments; (2) Adjustability—users can precisely control intervention strength and direction via a single parameter Durmus et al. (2024); (3) Composability—multiple steering methods can be combined for flexible control (Bayat et al., 2025). These properties enable efficient and fine-grained manipulation of model behaviors while enhancing interpretability. Particularly, recent works show that steering features extracted from SAEs (Huben et al., 2024; Templeton et al., 2024) are more interpretable and monosemantic, leading to better steering effects with fewer side effects  (Zhao et al., 2024; Farrell et al., 2024; Chalnev et al., 2024; Ferrando et al., 2024; Soo et al., 2025; Chalnev et al., 2024; Mayne et al., 2024).
@@
Early works note that neural networks may encode concepts linearly in activation spaces (Mikolov et al., 2013; Pennington et al., 2014), and the linear representation hypothesis has since evolved (Nanda et al., 2023; Park et al., 2024). Building on this, activation-based methods add scalable steering vectors to activations, enabling adjustability and composability. Prompt-based methods, such as many-shot prompt (Anil et al., 2024; Agarwal et al., 2024), achieve similar control via natural language. Decoding-based techniques modify generation logic in a similar manner (Dathathri et al., 2020; Yang and Klein, 2021).
@@ 3.1 Overview
  • Reasoning Pattern: constraining the length of reasoning processes, balancing parametric and contextual knowledge (Zhao et al., 2024), eliciting deliberate thinking, and enforcing discipline-specific reasoning structures Chen et al. (2025).
@@ 3.6 Evaluators Module
The evaluators module assesses the quality of outputs generated by a steered model by processing result files from evaluation datasets. Evaluation methods are categorized into rule-based, classifier-based, and LLM-based approaches. Given the diversity of steering concepts, our framework supports multiple evaluation dimensions and enables user-defined evaluations through an adaptive LLM-based strategy. Inspired by AXBENCH (Wu et al., 2025), we leverage powerful models (e.g., GPT-4) to handle a wide range of steering concepts. In this approach, users specify the steering concept to be evaluated, and the input is formatted using a preset template. Various evaluation metrics, including concept, instruction, and fluency scores, are then computed to measure steering effectiveness comprehensively.
3.7 Steering Methods supported in EasyEdit2
Prompt-based Steering.
This category, which encompasses manually designed prompts and auto-generated prompts methods  Wu et al. (2025), directly influences the model’s responses through prompt engineering.
Activation-based Interventions.
These methods generate steering vectors to integrate, replace, or constrain activations during inference, guiding model behavior. One of the core methods, Contrastive Activation Addition (CAA) Rimsky et al. (2024), steers language models by generating steering vectors, which compute activation differences between positive and negative example pairs. LM-Steer Han et al. (2024a) takes a different approach by applying a lightweight linear transformation to output embeddings. SAE Feature Steering leverages features extracted from SAEs, enabling users to select SAE features associated with specific concepts and apply them as steering vectors. The Steering Target Atoms (STA) method extends CAA by leveraging a Sparse Autoencoder (SAE) to refine its steering vectors.
Decoding-based Control.
@@ 4.1 Experimental Settings
In our experiments, we primarily evaluate our framework on safety and sentiment, and employ the Gemma-2-9B (Team et al., 2024) and the Qwen2.5-7B (Yang et al., 2024a) as the base models.
@@
Baseline 58.29 4.619 59.38 4.901 58.38 4.708 55.54 5.029
CAA 64.72 4.662 72.76 4.949 66.88 4.371 66.32 5.050
STA* 63.55 4.672 72.78 4.954
LM-Steer 63.8 4.422 60.38 4.147 73.47 4.425 59.38 3.320
PromptAuto 58.96 4.481 66.96 4.021 60.13 4.676 62.16 4.140
@@
Table 2: Performance comparison of steering methods on sentiment and safety tasks. DR denotes Defense Rate, FL indicates Fluency, and POS represents Positive Rate. The best results are highlighted in blue.
For sentiment, we similarly sample 2,000 instances from SST-2 (Socher et al., 2013). For evaluation, we use the Neutral dataset constructed by Han et al. (2024a) and apply HuggingFace’s sentiment classifier (Wolf et al., 2020) to assess the outputs. The sentiment score is the percentage of positive outputs.
We evaluate four steering methods: CAA, LM-Steer, STA, and PromptAuto, comparing their effectiveness across safety and sentiment dimensions. For CAA and STA, which require selecting model layers for intervention, we empirically select middle to late layers, using layer 24 for Gemma and layer 16 for Qwen. Full hyperparameter settings are available in our EasyEdit GitHub repository [^4] https://github.com/zjunlp/EasyEdit/tree/main/hparams/Steer/experiment_hparams.
The results largely met expectations, demonstrating that EasyEdit2 effectively has the ability to steer the behavior of LLMs.
4.2 Results
As shown in Table 2, all steering methods outperform the baseline. CAA and STA, which modify activations at inference time, achieve high defense rates and sentiment scores, demonstrating their effectiveness. LM-Steer shows improvements in some cases but is constrained by its need for additional parameter training, sensitivity to hyperparameters, and reliance on multi-label datasets. PromptAuto exhibits certain limitations, as its effectiveness depends heavily on prompt quality and the specific steering scenario.
Figure 4: A code snippet in EasyEdit2, where the CAA method shifts output language from English to Chinese.
@@ 5 Demonstration
Figure 5 displays our online demo built with Gradio, which is directly accessible via the web. The demo is organized into two tabs: one for test-time steering and one for SAE-based fine-grained control, where users can specify or search for SAE features to steer the model. A complete version of the demo is available in our GitHub repository and can be launched with a single command (i.e., python app.py).

### v2

@@ Abstract
In this paper, we introduce EasyEdit2, a framework designed to enable plug-and-play adjustability for controlling Large Language Model (LLM) behaviors. EasyEdit2 supports a wide range of test-time interventions, including safety, sentiment, personality, reasoning patterns, factuality, and language features. Unlike its predecessor, EasyEdit2 features a new architecture specifically designed for seamless model steering. It comprises key modules such as the steering vector generator and the steering vector applier, which enable automatic generation and application of steering vectors to influence the model’s behavior without modifying its parameters. One of the main advantages of EasyEdit2 is its ease of use—users do not need extensive technical knowledge. With just a single example, they can effectively guide and adjust the model’s responses, making precise control both accessible and efficient. Empirically, we report model steering performance across different LLMs, demonstrating the effectiveness of these techniques. We have released the source code on GitHub [^1] https://github.com/zjunlp/EasyEdit along with a demonstration notebook. In addition, we provide an online system [^2] http://easyedit.zjukg.cn/ for real-time model steering, and a demo video [^3] https://www.youtube.com/watch?v=AkfoiPfp5rQ for a quick introduction.
@@ 1 Introduction
To this end, we introduce EasyEdit2—a new, easy-to-use steering framework for editing LLMs. Building on the foundation of the legacy EasyEdit Yao et al. (2023); Wang et al. (2024b); Zhang et al. (2024), EasyEdit2 features an entirely new architecture designed to enhance plug-and-play capabilities and improve adjustability when steering LLMs. Currently, a variety of steering methods—including prompt-based steering, activation-based interventions (Turner et al., 2023; Rimsky et al., 2024; Wang et al., 2024c; Hartvigsen et al., 2023; Scialanga et al., 2025), decoding-based control—exist, yet they remain fragmented and require custom implementations and significant expertise. Thus, we develop the steering vector generator module and the steering vector applier module to automatically generate steering vectors and apply these vectors for intervention (if employing prompt-based steering, generating a steering vector is unnecessary). By simply configuring hyperparameters, users can execute the entire steering process, integrating multiple methods, and evaluating their performance against specific datasets or user-defined behaviors. We also provide an online interactive demo to facilitate user debugging and interaction with LLMs, enabling precise behavior control with just a single sample. To further assist users, our framework is released under the MIT License, ensuring open access and flexibility for use, modification, and distribution.
Unlike prior work, such as AXBENCH (Wu et al., 2025a), which designs data to evaluate steering methods across fine-grained concepts, and Dialz (Siddique et al., 2025), which focuses on the use and interpretability of activation-based steering vectors, EasyEdit2 provides a more flexible and user-friendly framework. Specifically, this framework enables users to combine multiple steering methods and merge steering vectors, improving single-objective steering and enabling multi-objective steering across diverse tasks. To achieve this, EasyEdit2 features a steering vector library for reusing existing vectors and supports algebraic merging, allowing the combination of distinct vectors without manual reengineering of the underlying model. Additionally, EasyEdit2 introduces few-shot steering, where a single contrastive example can guide effective vector generation, reducing data requirements for precise behavior control.
@@ 2 Background
Inference-time steering modifies model behavior during inference through prompt-based (Wu et al., 2025a), activation-based (Zou et al., 2023; Stolfo et al., 2024; Bartoszcze et al., 2025; Wehner et al., 2025; Wu et al., 2025b; Sun et al., 2025), and decoding-based methods (Liang et al., 2024). Compared to parameter fine-tuning methods Han et al. (2024b), inference-time intervention offers several key advantages: (1) Pluggability—steering methods can be seamlessly applied or removed without changing model weights, whether through activation modification, prompt-based guidance, or decoding adjustments; (2) Adjustability—users can precisely control intervention strength and direction via a single parameter Durmus et al. (2024); (3) Composability—multiple steering methods can be combined for flexible control (Bayat et al., 2025). These properties enable efficient and fine-grained control of model behaviors while enhancing interpretability. Particularly, recent works show that steering features extracted from SAEs (Huben et al., 2024; Templeton et al., 2024) are more interpretable and monosemantic, leading to better steering effects with fewer side effects  (Zhao et al., 2024; Farrell et al., 2024; Chalnev et al., 2024; Ferrando et al., 2024; Mayne et al., 2024; Soo et al., 2025).
@@
Early studies suggest neural networks may encode concepts linearly in activation space (Mikolov et al., 2013; Pennington et al., 2014), a view refined by recent work (Nanda et al., 2023; Park et al., 2024). Building on this, activation-based methods steer model behavior by adding scalable vectors to activations, enabling adjustable and composable control. Prompt-based methods (Anil et al., 2024; Agarwal et al., 2024) achieve similar control through natural language, while decoding-based methods (Dathathri et al., 2020; Yang and Klein, 2021) achieve control by altering decoding logic.
@@ 3.1 Overview
  • Reasoning Pattern: constraining the length of reasoning processes, balancing parametric and contextual knowledge (Zhao et al., 2024), eliciting more deliberate and structured thinking, and enforcing discipline-specific reasoning structures Chen et al. (2025).
@@ 3.3 Steering Vector Applier Module
3.4 Steering Vector Library and Merging
A key innovation in EasyEdit2 is developing a steering vector library with support for vector merging.
Steering Vector Library.
@@ 3.7 Evaluators Module
The evaluators module assesses the quality of outputs generated by a steered model by processing result files from diverse evaluation datasets. Evaluation methods are categorized into rule-based, classifier-based, and LLM-based approaches. Given the diversity of steering concepts, our framework supports multiple evaluation dimensions and enables flexible, user-defined evaluations through an adaptive LLM-based strategy. Inspired by AXBENCH (Wu et al., 2025a), we leverage powerful models (e.g., GPT-4) to handle a wide range of complex steering concepts. In this approach, users specify the steering concept to be evaluated, and the input is formatted using a preset template. Various evaluation metrics, including concept relevance, instruction relevance, and fluency scores, are then computed to comprehensively measure steering effectiveness.
@@ 4.1 Experimental Settings
In our experiments, we primarily evaluate our framework on safety and sentiment in Gemma-2-9B (Team et al., 2024) and the Qwen2.5-7B (Yang et al., 2024a) models. We consider two settings: single-task settings, where each method is trained and tested separately on individual tasks; and multi-task settings, where methods are trained and evaluated jointly across multiple tasks.
For safety, we evaluate on 1,200 prompts from RealToxicityPrompts (Gehman et al., 2020), with toxicity scores computed using the Perspective API [^4] https://perspectiveapi.com. For sentiment, we evaluate on the Neutral subset constructed by Han et al. (2024a), using a HuggingFace sentiment classifier (Wolf et al., 2020) to assess positivity. Full dataset and evaluation details are provided in Appendix B.1. Full hyperparameter configurations are available at our EasyEdit GitHub repository [^5] https://github.com/zjunlp/EasyEdit/tree/main/hparams/Steer/experiment_hparams.
In the single-task setting, we evaluate CAA, LM-Steer, STA, and Promptauto (details in Appendix A). For CAA and STA, we apply interventions at layer 24 for Gemma and layer 16 for Qwen.
To enable multi-task generalization, we further introduce a steering vector merging setup. Specifically, we merge CAA-derived vectors obtained from safety and sentiment tasks using Linear, TIES, or DARE-TIES (details are in Section 3.4), and evaluate the resulting vector jointly on both tasks—allowing a single intervention to influence multiple behavioral objectives.
@@
Single-task Steering
Baseline 58.30 4.618 58.80 4.901 58.30 4.684 55.54 5.029
CAA 64.80 4.661 72.76 4.949 66.89 4.370 66.32 5.050
STA* 63.64 4.671 72.78 4.954
LM-Steer 63.80 4.422 60.38 4.147 73.47 4.425 59.38 3.320
Promptauto 59.13 4.335 66.96 4.021 60.13 4.547 62.16 4.140
Multi-task Steering with Merged CAA Vectors (Safety + Sentiment)
Linear 67.47 4.694 75.38 4.982 73.81 4.375 71.84 4.745
TIES 68.06 4.706 76.44 4.982 72.73 4.389 69.56 4.745
DARE-TIES 68.06 4.719 75.68 5.013 72.14 4.388 70.96 4.729
@@
Table 2: Performance comparison of single-task and merged-vector steering methods. Single-task vectors are trained and tested separately on safety and sentiment, while merged CAA vectors are jointly evaluated on both. DR = Defense Rate, FL = Fluency, POS = Positive Rate. Best results are in bold, second-best are underlined.
4.2 Main Results
Activation-based methods such as CAA and STA are effective for safety and sentiment control in single-task settings.
Results in Table 2 show that CAA and STA consistently outperform other methods when trained and evaluated on individual tasks, benefiting from direct activation intervention. LM-Steer exhibits less stable performance due to its reliance on additional training and multi-label supervision, while Promptauto is sensitive to prompt formulation and task context.
Merged steering vectors demonstrate strong composability, enabling unified control across multiple objectives.
Table 2 shows that merging CAA-based vectors, separately trained on safety and sentiment tasks, using Linear, TIES, or DARE-TIES allows simultaneous control over both objectives. Notably, the merged vectors perform on par with—or even outperform—their single-task vectors, highlighting the efficiency and flexibility of multi-behavior steering.
Steering vectors enable precise and bidirectional adjustability via multipliers.
Figure 5 illustrates that applying positive or negative multipliers to steering vectors enables smooth and interpretable adjustment of safety and sentiment directions, validating the scalability and controllability of vector-based interventions.
Additional Experiments and Evaluation Details.
We further evaluate steering methods on the AXBENCH benchmark, which focuses on fine-grained concept control. EasyEdit2 has partially integrated AXBENCH evaluation, and the results (shown in Table 3 and detailed in Appendix B.3) indicate that prompt-based methods perform better in fine-grained scenarios, whereas activation-based methods are more effective for coarser, intensity-driven tasks. Further experimental details and analyses are provided in Appendix B.
5 Demonstration
@@
Figure 4 displays our online demo built with Gradio, which is directly accessible via the web. The demo is organized into two tabs: one for test-time steering and one for SAE-based fine-grained control (Appendix C.2), where users can specify or search for SAE features to steer the model. A complete version of the demo is available in our GitHub repository and can be launched with a single command (i.e., python app.py).
@@ Acknowledgements
Appendix A Steering Methods Supported by EasyEdit2
Prompt-based Steering.
This category, which encompasses manually designed prompts and auto-generated prompts methods  Wu et al. (2025a), directly influences the model’s responses through prompt engineering.
Activation-based Interventions.
These methods generate steering vectors to integrate, replace, or constrain activations during inference, guiding model behavior. One of the core methods, Contrastive Activation Addition (CAA) Rimsky et al. (2024), steers language models by generating steering vectors, which compute activation differences between positive and negative example pairs. LM-Steer Han et al. (2024a) takes a different approach by applying a lightweight linear transformation to output embeddings. SAE Feature Steering leverages features extracted from SAEs, enabling users to select SAE features associated with specific concepts and apply them as steering vectors. The Steering Target Atoms (STA) Wang et al. (2025) method extends CAA by leveraging a Sparse Autoencoder (SAE) to refine its steering vectors.
Decoding-based Control.
Appendix B Experimental Details
B.1 Safety and Sentiment Task Setup
For sentiment, we similarly sample 2,000 instances from SST-2 (Socher et al., 2013) as training data. For evaluation, we use the Neutral dataset constructed by Han et al. (2024a) and apply HuggingFace’s sentiment classifier (Wolf et al., 2020) to evaluate the outputs. The sentiment score is the percentage of positive outputs.
B.2 Adjustability of Steering Vectors
We investigate the adjustability of merged steering vectors on the Gemma-2-9b model by modulating their effect with different multipliers during inference. Using CAA-based vectors from safety and sentiment tasks, we test three merging strategies: Linear, TIES, and DARE-TIES. Details of these merging strategies are provided in Section 3.4. Each steering vector is scaled by a multiplier from $-2$ to $2$ and applied during inference.
As shown in Figure 5, scaling enables smooth, bidirectional control over safety (DR) and sentiment (POS): positive scaling enhances target behaviors, while negative scaling suppresses them. Notably, fluency (FL) also increases with the scaling factor, despite not being directly optimized, potentially due to increased model confidence or alignment with learned directions.
Figure 5: Adjustability analysis of steering vectors. Adjusting the magnitude (multiplier) of merged vectors enables smooth, bidirectional control over both safety (DR) and sentiment (POS), with fluency (FL) shown as an auxiliary measure. Bars show DS or POS metrics; lines show fluency (FL) as an auxiliary measure.
B.3 AxBench Evaluation Setup and Results
B.3.1 Datasets
We adopt the $\mathrm{D}^{9B}_{L20}$ subset from the CONCEPT500 dataset in AXBENCH, corresponding to the Gemma-2-9b-it model. As this subset is based on the 20th layer of the model, layer-specific steering methods like CAA and STA are configured to intervene at layer 20 accordingly. Following the methodology of Wu et al. (2025b), we employ the generated preference training data from this subset as the supervisory signals for steering. The dataset consists of pairs of input instructions and responses, with and without the targeted steering concept, enabling effective learning of steering vectors.
B.3.2 Evaluation
We conduct experiments on the Gemma-2-9b-it model in an instruction-following setup, where instructions are randomly sampled from Alpaca-Eval Li et al. (2023). The model generates responses while undergoing in-place forward pass interventions using the tested steering methods. To ensure comparability, we adopt the same prompt templates as those used in AXBENCH.
For each steering concept, we sample 10 instructions from Alpaca-Eval and generate corresponding responses. Outputs are then evaluated by GPT-4o-mini on discrete metrics scored in $\{0,1,2\}$:
  • Concept: How well the response expresses the intended concept.
  • Instruction: How well the response aligns with the given instruction.
  • Fluency: The linguistic quality and readability of the response.
  • Harmonic Mean (HM): The overall score combining the above three, penalizing poor performance in any single aspect.
B.3.3 Results
We evaluate several steering methods: Baseline, Promptauto, CAA, STA, and LM-Steer. Full hyperparameters are available in our GitHub [^10] https://github.com/zjunlp/EasyEdit/tree/main/hparams/Steer/experiment_hparams.
Table 3 summarizes the results on AXBENCH. Promptauto achieves the best performance, highlighting its advantage in handling fine-grained, concept-sensitive control tasks. Compared to its relatively weaker results on broader safety and sentiment tasks (Table 2), this suggests that prompt-based methods offer stronger generalization in nuanced settings, whereas activation-based methods such as CAA and STA are more effective for coarser, intensity-driven control.
Method Concept Instruct Fluency HM
Baseline 0.097 1.999 1.086 0.112
Promptauto 0.922 1.873 1.147 0.955
CAA 0.323 1.842 1.120 0.343
STA 0.382 1.551 0.734 0.199
LM-Steer 0.327 0.105 0.015 0.002
Table 3: Performance comparison of steering methods on Gemma-2-9B-it evaluated on AXBENCH using EasyEdit2. Best results are in bold, second-best are underlined.
Appendix C EasyEdit2 Usage Demonstration
We demonstrate the use of EasyEdit2 through two representative interfaces: a code snippet and an interactive online demo.
C.1 Code Snippet
Figure 6 illustrates the complete workflow of behavior steering in EasyEdit2 using the CAA method.
Figure 6: A code snippet in EasyEdit2 using CAA to steer model from neutral to emotionally expressive.
C.2 Online Demo
Figure 7 shows the SAE-based fine-grained control tab in online demo, where users search for features and adjust steering strength to modify outputs.
Figure 7: SAE-based fine-grained manipulation tab, showing feature-based steering for model control. In this example, the user inputs “chinese” to search for related SAE features. After selecting the relevant features, the steering strength is set, and the user provides the prompt “Tell me a story full of culture.” The resulting output is steered to emphasize Chinese culture, including food, art, and social elements, demonstrating the effectiveness of fine-grained steering in guiding the model’s response based on cultural context.